OpenStudy (anonymous):

Write equation in point slope, slope intercept, and standard. (-4,9) (6,-6)

OpenStudy (anonymous):

m = 15/ -10 = -3/2 Point-slope form: y+6 = ( -3/2)(x -6) Slope-intercept form: y +6 = -3x/2 +1/2 Standard form 2y +12 = -3x + 18 =>3x +2y = 6